a 60 kg marathon runner runs up the stairs to the top of Chicago’s 443m tall Willis Tower. to lift the top in 15 minutes, what must be her average power output? how many such girls can match the power of an average horse?
Answer #1

Mass of the runner, m= 60kg

Acceleration due to gravity, g = 9.8 m/s2

Time taken t= 15 min = 900s

Height to climb h= 443m

Now Power, P = Energy per unit time = E/t = mgh/t = (60*9.8*443) /900

So P = 289.43 Watt

Power required by average horse = P/1 horsepower = P/746 = 0.3879hp